# Next-Gen Software Engineering: AI-Assisted Big Models

### v1

@@ 4 AI in Software Engineering
Therefore, with respect to the AI-SEAL taxonomy Feldt et al. (2018), we consider the goals of AI application as missing and add them as a separate dimension with the three facets of understanding, generation and improvement. Furthermore, we refine the phases of application according to IEEE (2022) by extending the software production into the tasks of software engineering including requirements, architecture & design, construction, testing, and maintenance. For the overall SE process, we differentiate its management and its economics. We also extend software operations to software configuration and software execution. Finally, we limit the levels of autonomy in resembles of the levels of autonomous driving Barabas et al. (2017) into four levels of support from recommendations to full automation: (1) AI assistance, where the developer is in full control and receives recommendations to chose from, (2) AI-assisted selection where the AI preselects options, (3) AI-based partial automation where the AI selects options in simple, standard cases, and (4) AI-based full automation where the AI operates without the developer. As of today, level 1 is the most often used and level 4 is by far from becoming realistic. Whether there will be further differentiation in the preference for automation is also an open question.
This novel AI4SE taxonomy is shown in Figure 4. Due to space limitations, we will focus on the state of the art in applying ML to software development only. In addition, the fidelity and applicability of this new taxonomy is examined by presenting mainly recent and highly cited publications.
Figure 4: AI4SE - The Taxonomy of AI-assisted Software Engineering
Figure 5: AI-assisted Software Development
@@
Körner et al. (2014) contributes to requirements understanding, generation and improvement by presenting an AI-based automation approach to requirements engineering that begins by converting natural language into an Eclipse Modeling Framework (EMF) model. It then applies linguistic rules to identify errors, such as ambiguities or incorrect quantifiers, and provides suggestions for requirements analysts to make final decisions. This approach supports the entire requirements elicitation and change process.
Another requirements improvement is given in Perini et al. (2012) presenting the prioritisation of requirements by combining the preferences of project stakeholders with approximations of the order of requirements computed by ML techniques.
@@
For Architecture & design understanding and generation, Bhat et al. (2019) describes the automated curation of design decisions to support architectural decision-making. It helps software architects by organizing and recommending design decisions based on previous cases and contextual information of the current project. By leveraging existing design knowledge, the approach analyses historical data and design choices to improve the quality and consistency of architectural designs.
@@
The list of publications on AI-assisted coding is huge. Major developments are for example described in Gupta and Sundaresan (2018) for code understanding. It introduces DeepCodeReviewer that uses deep learning to recommend code reviews for common issues based on historical peer reviews. It assess the relevance of reviews to specific code snippets, suggests appropriate reviews from a repository of common feedback, and improves code reviews by focusing on defect detection. Guo et al. (2020) describes GraphCodeBERT, a pre-trained model for programming languages that incorporates data flow semantics rather than just code syntax. GraphCodeBERT demonstrates its performance both in code understanding for code search and clone detection, and in code generation and improvement through code translation and code refinement. Ceran et al. (2023) presents a study focused on predicting software quality using defect density as a key feature representing quality to achieve higher accuracy in software quality prediction compared to previous studies. The research shows that data pre-processing, feature extraction and the application of ML algorithms significantly improve prediction accuracy.
For code generation, Bird et al. (2023) discusses early experiences of developers using GitHub Copilot, which uses a language model trained on source code. Guided by Copilot, developers can write code faster than a human colleague, potentially accelerating development. Three empirical studies with Copilot highlight the different ways developers use Copilot, the challenges they face, the evolving role of code review, and the potential impact of pair programming with AI on software development. Svyatkovskiy et al. (2020) discusses IntelliCode Compose, a multilingual code completion tool that predicts entire sequences of code tokens up to full lines of code. The generative transformer model has been trained on 1.2 billion lines of Python, C#, JavaScript and TypeScript code.
Bader et al. (2019) presents Getafix, a tool for fixing common bugs by learning from previous human-written fixes, for code improvement. It uses hierarchical clustering to group bug fix patterns into a hierarchy from general to specific, and a ranking system based on the context of the code change to suggest the most appropriate fix. Another debugging approach, DeepDebug, is presented in Drain et al. (2021), which has been trained by mining GitHub repositories to detect and fix bugs in Java methods.
Since different software testing techniques are complementary, reveal different types of defects and test different aspects of a program, Lenz et al. (2013) presents for test understanding and improvement an ML-based approach to link test results from different techniques, to cluster test data based on functional similarities, and to generate classifiers according to test objectives, which can be used for test case selection and prioritisation.
@@
Tufano et al. (2022) discusses a test generation approach for writing unit test cases by generating assert statements. The approach uses a transformer model that was first pre-trained on an English text corpus, further semi-supervisedly trained on a large source code corpus, and finally fine-tuned for the task of generating assert statements for unit tests. The assert statements are accurate and increase test coverage.
For test improvement, Zhao et al. (2015) discusses test case prioritisation based on source code changes, software quality metrics, test coverage data, and code coverage-based clustering. It reduces the impact of similar test cases covering the same code and improves fault detection performance.
@@
For the understanding, generation and improvement of configurations for software deployments, Tatineni and Mustyala (2021) explores the role of ML in DevOps for intelligent release management. It suggests combining continuous monitoring, predictions of the likelihood of deployment failures, root cause analysis, and pipeline optimisation to reduce deployment failures and improve release management efficiency and software quality.
@@
Solutions which address several goals and tasks are associated with the SE process: For the improvement of the SE process, Spieker et al. (2017) introduces Retecs, a method for automatically learning test case selection and prioritisation in continuous integration, aimed at minimising the time between code commits and developer feedback on failed tests. Retecs uses reinforcement learning to select and prioritise test cases based on their execution time, previous execution history and failure rates. It effectively learns to prioritise error-prone test cases by following a reward function and analysing past CI cycles.
For the understanding of the SE process, Lin et al. (2021) presents the T-BERT framework for generating trace links between source code and natural language artefacts such as requirements or code issues. It demonstrates superior accuracy and efficiency for software traceability, especially in data-limited environments.
The presented selection of recent research demonstrates the feasibility of the added dimensions and facets of the AI4SE taxonomy compared to Feldt et al. (2018). It is our belief that it is only a matter of time before research results are presented in primary studies for the open aspects. For further reading, Wong et al. (2023); Allamanis et al. (2018) may be used.
5 AI-Assisted Big Models in Software Engineering
@@
In addition to the development of AI applications in SE, the availability of models for SE for MDSE has increased significantly. This is also due to the fact that in addition to the naturalness hypothesis of Allamanis et al. (2018), there is {quoting} the modelling hypothesis: Software is (also) a formalized communication between humans and computers; model corpora, software corpora, and natural language corpora have similar statistical properties; the properties of model corpora can be employed to develop more efficacious software engineering tools.
In their critical review, Hailpern and Tarr (2006) suggest that while MDSE may have potential in the context of large-scale, distributed industrial software development, it is not a guaranteed success. However, the advent of big code and AI has opened up new avenues for integrating AI applications, particularly ML, into existing SE models. As demonstrated by Hamilton et al. (2017), the intrinsic formal graph structures of SE models can be leveraged to facilitate the preparation of models for AI/ML applications. Conversely, there is a necessity for elevated abstraction levels to enhance the efficacy of AI-assisted automated coding Pudari and Ernst (2023).
Figure 6: AI-assisted Software Modelling
@@
When both factors are considered together, it becomes evident that there are significant opportunities for MDSE.
@@
    The application of AI methods on Big Models and for Big Models-
@@
    The shift towards the new paradigm of pair modelling in SE, which altogether will turn into the next generation of software engineering.
@@
Regarding opportunity (1), Stórrle et al. (2014) presented the first entities of the SEMI Software Engineering Models Index, a catalogue of model repositories, and invited further contributions to SEMI. Hebig et al. (2016) used another approach of mining GitHub for projects including Unified Modelling Language (UML) models, which could well be combined with the SEMI initiative. As a result of numerous efforts towards SE model collections, the Lindholmen dataset was developed Robles et al. (2017) and reviewed in Robles et al. (2023).
@@
In view of opportunity (2), extracting representational models from code is in fact the opposite of using models to generate code more efficiently. Since the design, specification, and maintenance of such models can be complex and time-consuming, various techniques have been developed to extract models from code and/or execution traces. These techniques include the extraction of all types of models including for example the extraction of information models, see e.g. Burson et al. (1990); Murphy and Notkin (1996), of structural models, see e.g. Kazman et al. (1998); Guo et al. (1999), and of behavioural models, see e.g. Corbett et al. (2000); Lo et al. (2009). The process of extracting models from code and/or execution traces offers the advantage of retrieving models that are up-to-date with the code/traces, but it also carries the potential disadvantage of mismatch with the model representation/abstraction requirements.
@@
With regard to opportunity (3), there are initial studies on Big Models such as Ho-Quang et al. (2017) which explores the increasing role of modelling, especially in safety-critical software development. It surveyed a range of projects utilising the UML and identified collaboration as the primary rationale for employing models. This is because models facilitate communication and planning for joint implementation efforts within teams, including those who are not directly involved in modelling or are new to the team.
@@
For opportunity (4), the application of AI methods to Big Models is elaborated in first studies like:
  • Shcherban et al. (2021); Mangaroliya and Patel (2020) supporting the classification of UML diagrams and contributing to model understanding.
  • Babur (2016) presenting an approach for comparing and merging model variants by incorporating techniques from information retrieval, natural language processing, and machine learning and contributing to model understanding and improvement.
  • Baki and Sahraoui (2016) describing an approach to learn model transformations from examples of source and target model pairs and being a contribution for model generation.
@@
Last but not least for opportunity (5), the pair modelling paradigm naturally evolves from the insight that pair programming (see e.g. Bipp et al. (2008)) with AI tools based on large code models as pair programmers and partners in software development evolved into a very supportive application of AI in software engineering (see e.g. Dakhel et al. (2023)).
@@
Based on Big Models and large SE models models for ML, i.e. large models for ML trained with SE models, pair modelling will emerge as a new model-driven software development technique in which a software engineer and an AI tool collaborate in software development. The engineer or tool in the role of the driver writes or improves software artefacts, including models, while the other, in the role of the observer, reviews each element of the software artefact as it is typed into an artefact. The one in the role of observer is also in the role of navigator, considering systemic and strategic aspects of software development: The navigator identifies potential improvements and possible upcoming problems that need to be addressed later on if they are not to be avoided. This allows the driver to focus on the development aspects without losing sight of the crosscutting and overarching aspects of software development. Like in pair programming in pair modelling, the observer is used as an assurance of the overall resulting software quality and as a guide to high quality software engineering. The two, the driver and observer/navigator, can switch roles. Indee, the exact interplay of driver and observer/navigator in pair modelling depends on the capabilities of the AI tool and the level of support according to the AI4SE taxonomy, it can provide to the software engineer being in either role.

### v2

@@ 4 Artificial Intelligence in Software Engineering
Figure 4: AI4SE - The AI4SE Taxonomy of AI-assisted Software Engineering
The AI4SE taxonomy was inspired by numerous discussions concerning the application of AI in the field of SE. In the absence of a taxonomy that addressed the various options and intricacies of SE and/or AI, a new taxonomy was developed, drawing inspiration from other surveys (see section 6). The dimensions of the first level of AI4SE are categorised as follows: (1) the purpose of using AI for software engineering, (2) the target of using AI within software engineering, (3) the type of AI used, and (4) the level of autonomy/automation achieved by the AI.
The purpose dimension has the three facets of (1) understanding to gain insight, (2) generating to derive (parts of), and (3) improving to advance SE artefacts and/or SE processing. An approach may address multiple purposes. If it is hybrid in this sense, all its purposes are made explicit, e.g. to support understanding only is denoted by ’1’, or to support both understanding and generation is denoted by ’1, 2’.
The AI4SE application targets consist of the main SE activities as described by SWEBOK [73] and DevOps [51]: (1) Software development includes requirements engineering, architecture and design, coding, testing, and maintenance. (2) Software operations includes deployment, execution, and monitoring. Together with the (3) SE process including management and adjustment, the three form the target of application dimension in the taxonomy. Each of the target options can be performed with the support of models, based on models or even driven by models, or without them. For the sake of comprehension of the taxonomy, the model option is optional for each application target: If models are used, i.e. the application target is MDSE, it is marked with "(m)". Thus, the application target "1.m" refers to software development with models, and "1" to one without. Or, in a more fine-grained analysis, "1.d.m" would be used for model-based testing and "1.d" for testing without models. An approach may address several application targets. If it is hybrid in this sense, all of its goals can be made explicit, e.g. support for coding only is denoted by "1.c", support for both coding and testing is denoted by "1.c, 1.d", or support for coding and model-based testing is denoted by "1.c, 1.d.m". For simplicity, "h" can also be used to indicate a hybrid approach of several targets.
The types of AI are divided into (1) symbolic AI, i.e. knowledge representation, rule-based systems, and logical reasoning, (2) statistical AI, i.e. probabilistic modelling and inference, Bayesian networks, Markov models, Gaussian processes, etc., (3) classical machine learning, i.e. supervised, unsupervised, and reinforcement learning, feature engineering and selection, etc., (4) deep learning, i.e. neural networks, transformers, transfer learning, self-supervised learning, attention mechanisms, etc., (5) generative AI (GenAI), i.e. large language models (LLM), generative adversarial networks, diffusion models, variational autoencoders, autoregressive models, contrastive learning, reinforcement learning from human feedback, etc., (6) agentic AI, i.e. multi-agent systems, planning and decision making, hierarchical reinforcement learning, etc., and (7) general purpose AI, i.e. meta-learning and world models. Today, all types (1) to (6) are used in software engineering, but not (7) due to its non-existence. An approach can also use a combination of AI types, i.e. it can be hybrid. This can be made explicit, e.g. by denoting "1, 2" for a hybrid approach that uses both symbolic and classical machine learning. For simplicity, "h" can just be used to denote a hybrid approach that uses more than one type of AI.
We define the levels of autonomy/automation similarly to the levels of autonomous driving [9] as five levels from no support to recommendations to full automation: (1) none, (2) AI-assisted options, where the developer is in full control and receives recommendations to choose from, (3) AI-assisted selection, where the AI pre-selects options, (4) AI-based partial automation, where the AI selects options in simple, standard cases, and (5) AI-based full automation, where the AI operates without the developer. So far, level 3 is the most common, level 4 is on the rise and level 5 is rather unrealistic for complex, industrial-scale software. Higher levels of autonomy typically include the capabilities of lower levels of autonomy; when presenting the level of autonomy of an approach, the highest level of autonomy is given.
Paper Summary Purpose Target Type Level
[55] T-BERT generates trace links between source code and natural language artefacts. It represents a significant advancement in automated software traceability, offering improved accuracy, efficiency, and practicality. 2, 3 1 4 4
[49] NLP approach for improving requirements engineering in the automotive industry: It identifies flaws, such as ambiguous words, incomplete process descriptions, or incorrect quantifiers. 3 1.a 1 3
[63] CBRank method for prioritising software requirements: It uses stakeholder input and requirement attributes to generate an optimal ranking and enables adaptive elicitation by refining the prioritisation based on real-time feedback. 3 1.a, 2.a, 3.1 1, 3 3
[12] ADeX is a tool designed to automate the curation of architectural design decisions in software development. It helps automate the management of architectural decisions, reducing documentation effort while improving decision accuracy. 1, 3 1.b, 2.a 1, 3 4
[14] Explores the early experiences of developers using the AI-augmented coding assistant GitHub Copilot. It may may expand to debugging, code review, and software maintenance, requiring better trust and provenance tracking. 1, 2, 3 1.c 5 4
[78] IntelliCode Compose is a code completion tool that uses a generative transformer model to predict sequences of code tokens, including entire lines of code. It has been trained on a massive dataset of 1.2 billion lines of code across multiple programming languages. 2, 3 1.c 2, 4, 5 4
[6] Getafix is an automated bug-fixing tool that learns from past, human-written fixes to suggest human-like fixes for static analysis warnings. 2, 3 1.c, 1.d 2, 3 3
[38] GraphCodeBERT improves code understanding by incorporating dataflow graphs into the BERT transformer architecture. Data flow graphs represent dependencies between code variables and improve semantic understanding over treating code as a simple sequence of tokens. 1, 3 1.c, 1.d, 1.e 2, 4, 5 4
[40] DeepCodeReviewer is an automatic code analysis system that learns from historical peer reviews to identify and suggest relevant reviews for code snippets. 1, 3 1.c, 1.d, 3.a 2, 3, 4 4
[28] Introduces DeepDebug or automated bug detection and fixing in Java: It generates non-deletion fixes that modify instead of just deleting code 3 1.c, 1.e 4, 5 4
[81] Approach improves code review automation with DL by working with pre-trained models on raw source code and English text. 3 1.c, 1.d 5 3
[79] The book discusses the application of AI beside others for infrastructure monitoring and anomaly detection, predictive maintenance, and build pipeline optimization. 1, 3 1.c, 1.d, 2.a, 3.b 2, 3 4
[16] Presents RepairAgent for automated program repair that allows to autonomously plan and execute actions, such as gathering information, searching for code snippets, and validating fixes, by interacting with a set of specifically designed tools. 3 1.c, 2.a 5, 6 4
[52] Approach links different types of testing information to help automate the generation of equivalence classes for testing. The linking supports tasks such as test reduction, prioritisation and selection. 1, 2, 3 1.d 2, 3 3
[87] Presents a hybrid approach to test case prioritization for regression testing aiming to minimise testing costs. Aims to enhance software testing efficiency by optimising test case execution order. 3 1.d, 1.e 1, 2, 3 3
[75] Introduces Retecs for automatic test case prioritisation and selection in continuous integration. It uses test case meta-data to adapt to changes in the environment and prioritise error-prone test cases. 3 1.d, 2.a 3, 4 4
[21] Explores machine learning-based ensemble methods to predict software quality more accurately than previous approaches 1, 3 1.d, 3.a 3 4
Table 1: AI-assisted software engineering (sorted by application targets)
This novel AI4SE taxonomy [^2] If you want to work on it further, please contact the author for the ontology file which further refines the targets of applications and the AI types. is shown in Figure 4. The fidelity of this new taxonomy AI4SE is verified by reviewing and classifying highly cited publications, but also recent publications to demonstrate the applicability of this taxonomy. The classification of the papers is given in table 1. Additional details on the papers that target mainly the software development are given here:
@@
For requirements, [49] presents an AI-based automation approach to requirements engineering that begins by converting natural language into an Eclipse Modelling Framework (EMF) model. It then applies linguistic rules to identify errors, such as ambiguities or incorrect quantifiers, and provides suggestions for requirements analysts to make final decisions. This approach supports the entire requirements elicitation and change process. Another requirements improvement is given in [63] presenting the prioritisation of requirements by combining the preferences of project stakeholders with approximations of the order of requirements computed by ML techniques.
@@
For architecture & design, [12] describes the automated curation of design decisions to support architectural decision-making. It helps software architects by organizing and recommending design decisions based on previous cases and contextual information of the current project. By leveraging existing design knowledge, the approach analyses historical data and design choices to improve the quality and consistency of architectural designs.
@@
The list of publications on AI-assisted coding, testing, and maintenance is huge. Major developments are for example described in [40] for code understanding. It introduces DeepCodeReviewer that uses deep learning to recommend code reviews for common issues based on historical peer reviews. It assess the relevance of reviews to specific code snippets, suggests appropriate reviews from a repository of common feedback, and improves code reviews by focusing on defect detection. [38] describes GraphCodeBERT, a pre-trained model for programming languages that incorporates data flow semantics rather than just code syntax. GraphCodeBERT demonstrates its performance both in code understanding for code search and clone detection, and in code generation and improvement through code translation and code refinement. [21] presents a study focused on predicting software quality using defect density as a key feature representing quality to achieve higher accuracy in software quality prediction compared to previous studies. The research shows that data pre-processing, feature extraction and the application of ML algorithms significantly improve prediction accuracy. [14] discusses early experiences of developers using GitHub Copilot, which uses a language model trained on source code. Guided by Copilot, developers can write code faster than a human colleague, potentially accelerating development. Three empirical studies with Copilot highlight the different ways developers use Copilot, the challenges they face, the evolving role of code review, and the potential impact of pair programming with AI on software development. [78] discusses IntelliCode Compose, a multilingual code completion tool that predicts entire sequences of code tokens up to full lines of code. The generative transformer model has been trained on 1.2 billion lines of Python, C#, JavaScript and TypeScript code. [6] presents Getafix, a tool for fixing common bugs by learning from previous human-written fixes. It uses hierarchical clustering to group bug fix patterns into a hierarchy from general to specific, and a ranking system based on the context of the code change to suggest the most appropriate fix. Another debugging approach, DeepDebug, is presented in [28], which has been trained by mining GitHub repositories to detect and fix bugs in Java methods.
@@
Since different software testing techniques are complementary, reveal different types of defects and test different aspects of a program, [52] presents an ML-based approach to link test results from different techniques, to cluster test data based on functional similarities, and to generate classifiers according to test objectives, which can be used for test case selection and prioritisation. [81] discusses a test generation approach for writing unit test cases by generating assert statements. The approach uses a transformer model that was first pre-trained on an English text corpus, further semi-supervisedly trained on a large source code corpus, and finally fine-tuned for the task of generating assert statements for unit tests. The assert statements are accurate and increase test coverage. [87] discusses test case prioritisation based on source code changes, software quality metrics, test coverage data, and code coverage-based clustering. It reduces the impact of similar test cases covering the same code and improves fault detection performance.
@@
For deployments, [79] explores the role of ML in DevOps for intelligent release management. It suggests besides other things combining continuous monitoring, predictions of the likelihood of deployment failures, root cause analysis, and pipeline optimisation to reduce deployment failures and improve release management efficiency and software quality.
@@
For the SE process, [75] introduces Retecs, a method for automatically learning test case selection and prioritisation in continuous integration, aimed at minimising the time between code commits and developer feedback on failed tests. Retecs uses reinforcement learning to select and prioritise test cases based on their execution time, previous execution history and failure rates. It effectively learns to prioritise error-prone test cases by following a reward function and analysing past CI cycles. [55] presents the T-BERT framework for generating trace links between source code and natural language artefacts such as requirements or code issues. It demonstrates superior accuracy and efficiency for software traceability, especially in data-limited environments.
Overall, the majority of the presented selection of recent research focuses on improving the software engineering process, either by improving coding, requirements engineering, testing, or the overall processes. A smaller number of approaches aim to improve the understanding of software engineering artefacts. There are also methods that focus on generating code or code-related content. A mixture of AI types are used, including symbolic, statistical, ML, DL and GenAI, as well as hybrid approaches. Many of the AI approaches offer AI-assisted selection, i.e. they provide suggestions, recommendations and rankings that can help developers, but require a developer to make the final decision on what to do. AI-based partial automation is also a common level, where tools automate part of the process, but may require a human to review or further integrate the results. The facets of the AI4SE taxonomy are adequately addressed, although not all options are covered. Nevertheless, it is only a matter of time before research results are presented in primary studies for the open combinations of the facets, as long as they provide advances beyond the state of the art.
5 AI-Assisted Model-Driven Software Engineering
The substantial increase in the utilisation of AI in SE is in parallel accompanied by a notable rise in the application of AI to MDSE. This phenomenon can also be explained by our novel modelling hypothesis, which is an extension of the naturalness hypothesis [3] to modelling artefacts in software engineering:
The modelling hypothesis
@@ 5 AI-Assisted Model-Driven Software Engineering
Software is (also) a formalized communication between humans and computers; model corpora, software corpora, and natural language corpora have similar statistical properties; the properties of model corpora can be employed to develop more efficacious software engineering tools.
Indeed, the advent of big code and AI has opened up new avenues for integrating AI into model-driven approaches, as posited by the novel modelling hypothesis. As demonstrated by [42], the intrinsic formal graph structures of SE models can also be leveraged to facilitate the preparation of models for gaining AI advantages in MDSE. This is analogous to the considerations in [64] on the necessity to elevate abstraction levels in language models to enhance the efficacy of AI-assisted automated coding.
Paper Summary Purpose Target Type Level
[50] Discusses NLP-driven model generation for more efficient model construction. Presents "purposive" meta-models that guide the interactions between domain experts and GenAI to generate effective prompts for model construction. 2 1.a.m, 1.b.m 1, 5 4
[22] Discusses the NLP-driven domain modelling by evaluating different LLMs and various prompt engineering techniques. 2 1.a.m, 1.b.m 5 4
[80] Details the creation and validation of a high quality dataset of UML diagrams. Provides a valuable research resource in the form of the UML dataset and presents a new, improved approach to UML diagram classification. 3 1.a.m, 1.b.m, 1.c,m, 1.d.m 4 3
[72] Presents a deep learning approach for automatically classifying images of UML diagrams. 1 1.b.m 4 3
[8] Presents an approach to automatically generate model transformations from examples. The approach is demonstrated through experimental evaluation on seven diverse transformation problems. 1, 2 1.b.m 1, 2, 3 4
[57] Explores the automated classification of UML class diagrams into forward-engineered and reverse-engineered diagrams. It provides a dataset with ground truth for the classification of class diagrams. 1, 3 1.b.m 3 3
[1] Presents an experience report detailing the use of LLMs to analyse and transform software variants represented in various formalisms (Java, UML, GraphML, statecharts). 1, 2 1.b.m, 1.c.m 5, 6 3
[29] Presents a software architecture for the AI-augmented automation of DevOps pipelines to address both architectural and technical challenges associated with the development and operation of complex systems. 1 1.b.m, 1.e.m 4 3
[5] Proposes a approach for statistically analysing large collections of models within the field of MDSE. The approach demonstrates its effectiveness in domain analysis and repository management, and shows the potential for applications such as model clone detection. 1 1.b.m, 1.e.m 2, 3 2
[54] Introduces FlowGen, a code generation framework that leverages multiple LLM agents to emulate different software process models (Waterfall, TDD, and Scrum). On standard benchmarks, the Scrum-based FlowGenScrum consistently outperforms other models and baselines. 2 1.c.m, 1.d.m, 1.e.m, 3.1.m, 3.2.m 5, 6 4
Table 2: AI-augmented model-driven software engineering (sorted by application targets)
As shown in Table 2, the growing trend in the adoption of AI for MDSE is reflected by various new research results:
Foundational work on ground truths for developing and benchmarking AI applications in MDSE is discussed in [80, 72] and [57]. [80] presents the creation and validation of a new, high-quality ground truth dataset of UML diagrams. The curated dataset of 2626 unique and accurately labelled images across six UML diagram types goes beyond the Lindholmen dataset of modelling artefacts [66, 67]. The new ground truth dataset was also presented for an improved approach to UML diagram classification. Also [72] presents a new dataset of images, manually labelled into four UML diagram types (class, activity, sequence, use case) and a non-UML category, which is used to develop new DL approaches with higher precision to automate UML diagram classification. Another ground truth dataset is described in [57]: A manually labelled subset of the Lindholmen dataset is used to explore the automated classification of UML class diagrams into forward-engineered (FwCD) and reverse-engineered (RECD) diagrams using supervised machine learning. The methodology in [5] uses techniques from information retrieval, NLP, and ML to enable an efficient comparison, clustering and clone detection of model artefacts.
[50] presents a methodology for AI-augmented MDSE as a continuous cycle of AI-generated model proposals by LLMs that are then reviewed and refined by humans. It seeks to bridge the gap between the natural language used by domain experts and the formal languages used in MDSE.
[22] investigates the feasibility of using LLMs to automate domain modelling in SE. It applies various prompt engineering techniques to a dataset of diverse domain modelling examples and finds that while LLMs show promising domain understanding and achieve relatively high precision in generating classes and attributes, their recall remains low, particularly for relationships. It concludes that full automation is currently impractical.
[8] presents a novel method for automatically generating model transformations from examples. It employs a three-step process: (1) analysing example model pairs to identify groups of similar transformations, (2) using genetic programming to learn basic rules for each group, and (3) refining these rules using simulated annealing to handle complex dependencies and value derivations. The method presented improves the efficiency and accuracy of learning complex model transformations.
[1] explores the use of LLMs to aid in the complex process of re-engineering existing software variants into a software product line (SPL) by identifying features to represent variability and commonalities. It describes the methodology used to prompt the LLMs, the results obtained, and the limitations encountered. There is a potential in assisting domain analysts and developers in transforming software variants into SPLs, but LLMs are limited in managing large variants or complex structures.
[29] presents a novel software architecture designed to address the increasing complexity of modern system development within a DevOps framework. The architecture integrates MDSE and AI techniques to enhance the entire software lifecycle, from requirements engineering to monitoring and maintenance. The architecture is evaluated through multiple industrial case studies.
[54] uses agentic AI where each agent embodies a specific development role such as requirement engineer, developer, or tester and collaborates via chain-of-thought prompting. The set of LLM agents emulate different software process models and are evaluated against four benchmarks (HumanEval, HumanEval-ET, MBPP, and MBPP-ET). The impact of individual development activities on code quality is also investigated, revealing the critical role of testing and the beneficial effects of design and code review.
All papers share the common view that the strengths of AI and MDSE should be combined to further improve the quality of software products and/or processes. While acknowledging the potential of AI-supported MDSE, the papers also highlight the limitations of LLMs, their sensitivity to prompt variations, their limited ability to handle complex inputs, and the risk of inaccuracies. There is a consensus that a fully automated approach is not yet feasible and that human review remains necessary to counteract inaccuracies introduced by AI. UML is seen in these papers as a common modelling language, both for creating models and as an input to AI models. There is also an expectation that AI techniques will support better model quality and improve software reuse by identifying and adapting components from previous projects. This will be supported by the use of model matching and clone detection techniques.
@@
So when AI and MDSE are combined, it becomes clear that there are significant opportunities:
@@
    The application of AI methods on Big Models and for Big Models.
@@
    The shift towards the new paradigm of pair modelling in SE, which altogether will turn SE into the next generation of software engineering.
@@
Regarding opportunity (1), [77] presented the first entities of the SEMI Software Engineering Models Index, a catalogue of model repositories, and invited further contributions to SEMI. [43] used another approach of mining GitHub for projects including UML models, which could well be combined with the SEMI initiative. As a result of numerous efforts towards SE model collections, the Lindholmen dataset was developed [66] and reviewed in [67]. For recent work on ground truth datasets see table 2.
@@
In view of opportunity (2), extracting representational models from code is in fact the opposite of using models to generate code more efficiently. Since the design, specification, and maintenance of such models can be complex and time-consuming, various techniques have been developed to extract models from code and/or execution traces. These techniques include the extraction of all types of models including for example the extraction of information models, see e.g. [20, 60], of structural models, see e.g. [48, 39], and of behavioural models, see e.g. [23, 56]. The process of extracting models from code and/or execution traces offers the advantage of retrieving models that are up-to-date with the code/traces, but it also carries the potential disadvantage of mismatch with the model representation and abstraction requirements.
@@
With regard to opportunity (3), there are initial studies on Big Models such as [44] which explores the increasing role of modelling, especially in safety-critical software development. It surveyed a range of projects utilising the UML and identified collaboration as the primary rationale for employing models. This is because models facilitate communication and planning for joint implementation efforts within teams, including those who are not directly involved in modelling or are new to the team.
@@
For opportunity (4), the application of AI methods to Big Models is elaborated in first studies as elaborated in table 2 above.
@@
Last but not least for opportunity (5), the pair modelling paradigm naturally evolves from the pair programming paradigm (see e.g. [13]), which has evolved into one of the killer applications of AI in software engineering, where AI-powered tools/agents act as pair programmers and partners in software development (see e.g. [25]).
@@
Based on Big Models and large SE models models for ML, i.e. large models for ML trained with SE models, pair modelling will emerge as a new model-driven software development technique in which a software engineer and an AI tool/agent collaborate in software development. The engineer or tool/agent in the role of the driver writes or improves software artefacts, including models, while the other, in the role of the observer, reviews each element of the software artefact as it is typed into an artefact. The one in the role of observer is also in the role of navigator, considering systemic and strategic aspects of software development: The navigator identifies potential improvements and possible upcoming problems that need to be addressed later on if they are not to be avoided. This allows the driver to focus on the development aspects without losing sight of the cross-cutting and overarching aspects of software development. Like in pair programming in pair modelling, the observer is used as an assurance of the overall resulting software quality and as a guide to high quality software engineering. The two, the driver and observer/navigator, can switch roles. Indeed, the exact interplay of driver and observer/navigator in pair modelling depends on the capabilities of the AI tool and the level of support according to the AI4SE taxonomy, it can provide to the software engineer being in either role.